2.1 Suppose you long one Australian dollar call and one Australian dollar put with an exercise exchange rate of 0.75 (USD/AUD). The price of call and the price of put is USD0.06. Using the 15-period (spot) exchange rates given below, answer parts (a)-(c). Time Spot Exchange Rate (USD/AUD) 0 0.35 1 0.40 2 0.45 3 0.50 4 0.55 5 0.60 6 0.65 7 0.70 8 0.75 9 0.80 10 0.85 11 0.90 12 0.95 13 1.00 14 1.05 15 1.10 (a) Compute the net call payo, net put payo and net combined payo. (b) Plot separately the Net Long call payo, the net long put payo and the net combined payo. (c) Name and provide denition of the shape of the combined payo obtained from part (b)? 


1
Expert's answer
2021-06-08T12:35:31-0400

(a) Computation of the net call pay, net put pay and net combined pay:-

Given Information:-

Call Strike Price(KC) = 0.75(US/AUD)

Put Strike Price(KP) = 0.65(US/AUD)

Call Premium = USD 0.04

Put Optrion = USD 0.08

Call Payoff=Max(0,Spot Price - Strike Price)

Put Payoff =Max(0,Strike Price - Spot Price)



(b)Plot separately the Net Long call pay, the net long put pay and the net combined pay

Call Payoff Plot:-




Put Payoff Plot:-



Combine Payoff Plot:-



(c)The combined option is called Strangle option Strategy.

It is the strategy in which investors take long position in Call Option and Put Option with different Strike price on the same underlying as well as same maturity.